Find the integral - х 2x dx OA. 12 + *2x-31.0 2x-3)? D. es 38.2-3.. durare HABER
Answer #1

n.de Sonnicos I = 22-3 lef. 2x-3= u so that; a= 4+3 8 2.dx=du ? da= du du ::D=5472) and ut du dut * I- Ź Ismo Show.du] = [ frif satisfied with the explanation, please rate it up..